OpenStudy (anonymous):

The formula for the surface area of a cylinder is A=2pr(r+h).Find h if A=80p and r=5. The formula for the surface area of a cylinder is A=2pr(r+h).Find h if A=80p and r=5. @Mathematics

OpenStudy (mathmagician):

\[80p=2p*r(r+h)\] \[40=r^2+r*h\] \[40-5^2=5*h\] h=3